# Towards Simple Machine Learning Baselines for GNSS RFI Detection

### v1

@@ Abstract
Machine learning research in GNSS radio frequency interference (RFI) detection often lacks a proper justification for the decisions made in deep learning-based model architectures. Our paper challenges the status quo in machine learning approaches for GNSS RFI detection, revealing the potentially misleading track of current research and highlighting alternative directions. Our position advocates for a shift in focus from solely pursuing novel model designs to critically evaluating the utility of complex black box deep learning methods against simpler and more interpretable machine learning baselines. Our findings demonstrate the need for the creation of simple baselines and suggest the need for more exploration and development of simple and interpretable machine learning methods for the detection of GNSS RFIs. The increment of model complexity in the state-of-the-art deep learning-based models often provides very little improvement. Thanks to a unique dataset from Swiss Air Force and Swiss Air-Rescue (Rega), preprocessed by Swiss Air Navigation Services Ltd. (Skyguide), we demonstrate the effectiveness of a simple machine learning baseline for GNSS RFI detection on real-world large-scale aircraft data containing flight recordings impacted by real jamming. The experimental results indicate that our solution successfully detects potential GNSS RFI with 91% accuracy outperforming state-of-the-art deep learning architectures. We believe that our work offers insights and suggestions for the field to move forward.
@@ 1 INTRODUCTION
Given the complexity of the GNSS RFI detection problem and motivated from the successes in other areas, such as natural language processing, computer vision and speech processing, state of the art deep learning architectures have been developed and applied to it. Although it is well intended, this line of research has rarely provided evidence of the necessity of deep learning. The state of the art deep learning approaches proceeded to introduce models of increased complexity with limited availability of training data.
The goal of this paper is to guide the research community towards more meaningful progress through a focus on studying the utility of their models by drawing useful but simple and interpretable machine learning baselines. We achieve this by introducing simple and effective baseline machine learning methods for near real-time detection of GNSS RFI from large amounts of aircraft onboard data collected from GPS and Attitude Heading Reference System (AHRS). We first train a machine learning model from historical aircraft onboard data, learning from normal data and known anomalies, and then apply this model to detect potential GNSS RFI in new, unseen flight recordings. We demonstrate that it performs on par or better than state of the art deep learning methods, thus challenging the efficiency and effectiveness of increasing model complexity to solve GNSS RFI detection problems.
@@ 3 Simple Machine Learning Baseline for GNSS RFI Detection
Among the numerous machine learning approaches for GNSS RFI detection presented in the past, there is often something consistent - they tend to overlook simpler baselines in pursuit of novelty. This leads to overly complex engineered solutions without much utility and a good rationale. Towards this end, we propose a simple method that exceeds the performance of state of the art deep learning approaches. As a result, this baseline helps us to understand the complexity of the underlying problem and provide a solid foundation for further investigation. Of note, our contribution is properly setting up known technical methods, proposing useful and explainable features for GNSS RFI detection and creating a strong baseline.
@@ 4.6 Discussion
Our findings indicate that a range of advanced deep learning models applied to GNSS RFI detection were consistently outperformed by a basic machine learning baseline. This outcome may be attributed to factors such as overfitting to normal data or the presence of high inherent randomness, which complicates the distinction between challenging anomalies and normal patterns. Ultimately, these deep learning techniques, while potentially effective for other time-series applications, offer limited advantages for GNSS RFI detection. Their added complexity and reduced interpretability are not warranted in this context. More concerning is the false sense of progress these methods can create, especially when evaluations against simpler baselines are insufficient or poorly executed.
In addition to highlighting the shortcomings of current approaches, we introduce a straightforward benchmark that can serve as a strong foundation for renewed research in GNSS RFI detection. We believe that such methods can help distinguish between anomalies that are readily identifiable and those that pose greater detection challenges—an essential step for driving meaningful progress. This perspective is supported by the notable consistency across all evaluated methods in terms of which anomalies were detected and which were missed.

### v2

@@ Abstract
Machine learning research in GNSS radio frequency interference (RFI) detection often lacks a clear empirical justification for the choice of deep learning architectures over simpler machine learning approaches. In this work, we argue for a change in research direction-from developing ever more complex deep learning models to carefully assessing their real-world effectiveness in comparison to interpretable and lightweight machine learning baselines. Our findings reveal that state-of-the-art deep learning models frequently fail to outperform simple, well-engineered machine learning methods in the context of GNSS RFI detection. Leveraging a unique large-scale dataset collected by the Swiss Air Force and Swiss Air-Rescue (Rega), and preprocessed by Swiss Air Navigation Services Ltd. (Skyguide), we demonstrate that a simple baseline model achieves 91% accuracy in detecting GNSS RFI, outperforming more complex deep learning counterparts. These results highlight the effectiveness of pragmatic solutions and offer valuable insights to guide future research in this critical application domain.
@@ 1 INTRODUCTION
Detecting GNSS RFI has increasingly attracted the attention of researchers applying deep learning methods, inspired by their success in domains like computer vision. While these models are capable of capturing complex temporal dependencies, their application is often motivated more by architectural and overall solution novelty than by demonstrable performance gains over classical, interpretable machine learning approaches. This reflects a broader concern in time series anomaly detection, where state-of-the-art deep models are frequently introduced without rigorous empirical validation, and often fail to significantly outperform robust, well-understood baseline methods Sarfraz, M.S., Chen, M.Y., Layer, L., Peng, K. and Koulakis, M. (2024).
This work emphasizes the practical value of simplicity in GNSS RFI detection by demonstrating that well-designed, interpretable machine learning baselines can rival or exceed the performance of more complex deep learning models. Leveraging large-scale flight data collected from GPS and Attitude Heading Reference Systems (AHRS), we develop and evaluate efficient models capable of detecting interference events in near real-time. Our approach involves training on historical flight data containing both normal and anomalous conditions, and applying the resulting models to new, unseen recordings. The results challenge the prevailing assumption that increasing model complexity is necessary for effective RFI detection, showing instead that thoughtfully engineered baselines can offer both robustness and transparency in operational environments — a view consistent with recent critical evaluations in time series anomaly detection literature Sarfraz, M.S., Chen, M.Y., Layer, L., Peng, K. and Koulakis, M. (2024).
@@ 3 Simple Machine Learning Baseline for GNSS RFI Detection
Recent approaches to GNSS RFI detection often emphasize algorithmic innovation, frequently adopting complex deep learning architectures. However, such complexity is not always matched by demonstrable gains in performance. Inspired by critical perspectives in time series anomaly detection research (e.g., Sarfraz, M.S., Chen, M.Y., Layer, L., Peng, K. and Koulakis, M. (2024)), we explore an alternative path: emphasizing simplicity, domain-aligned feature design, and robust evaluation. We present a lightweight machine learning baseline that surpasses a number of contemporary deep learning methods in identifying GNSS RFI. Importantly, this baseline also provides insights into the structure of the detection task itself. Our contribution lies in rigorously applying established techniques, crafting interpretable features tailored to GNSS behavior, and highlighting that well-grounded models can yield both performance and understanding.
@@ 4.6 Discussion
Recent critiques in the time series anomaly detection literature have highlighted the limitations of deep learning approaches when benchmarked against simple yet effective baselines Sarfraz, M.S., Chen, M.Y., Layer, L., Peng, K. and Koulakis, M. (2024). In alignment with these findings, our results show that a range of advanced deep learning models applied to GNSS RFI detection were consistently outperformed by a straightforward machine learning baseline. This underperformance may be attributed to factors such as limited training data and overfitting to normal flight behavior. While deep learning methods have demonstrated strong performance in many domains, their added complexity and reduced interpretability offer little advantage in this safety-critical setting. Perhaps more concerning is that the widespread use of these type of models—without proper comparison to simple baselines—can give a misleading impression of progress. To address this, we propose a transparent and easily reproducible benchmark that establishes a reliable foundation for future research in GNSS RFI detection. This method offers a practical means of distinguishing between trivial and genuinely difficult anomalies—an essential step toward advancing robust and meaningful solutions. Our conclusions echo broader insights in unsupervised anomaly detection, where simple models have repeatedly rivaled or exceeded the performance of state-of-the-art deep learning methods Sarfraz, M.S., Chen, M.Y., Layer, L., Peng, K. and Koulakis, M. (2024).